OpenStudy (anonymous):

Find three consecutive positive even integers such that the product of the second and third is twenty more than ten times the first integer

OpenStudy (anonymous):

Let n be the middle number. Then the numbers are: {n - 2, n, n + 2} Solve the following for n: n (n + 2) = 20 + 10 (n - 2) n = 8 {6, 8, 10}